OpenStudy (scarlett_feathers):

can someone help with this question? Will fan and medal. The perimeter of a rectangular garden is 350 feet. If the width of the garden is 82 feet, what is its length?

Vocaloid (vocaloid):

hint: 2L + 2W = 350 plug in W and solve for L

Vocaloid (vocaloid):

let me know if you have any additional questions

OpenStudy (scarlett_feathers):

\[2L+2(82)=350\] \[2L+164=350\] \[\frac{ 2l }{ 2}= \frac{ 186 }{ 2 }\] \[L=93\] @Vocaloid is this correct

Vocaloid (vocaloid):

yes, that's correct!

OpenStudy (scarlett_feathers):

can I tag you in another question?